Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by talaangoshtari » Mon Jul 27, 2015 4:02 am
A bag contains 3 white, 4 black, and 2 red marbles. Two marbles are drawn from the bag. What is the probability that the second ball drawn will be red if replacement is NOT allowed?

A. 1/36
B. 1/12
C. 7/36
D. 2/9
E. 7/9
Join the discussion
Source: — Problem Solving |

by GMATGuruNY » Mon Jul 27, 2015 4:13 am
talaangoshtari wrote:A bag contains 3 white, 4 black, and 2 red marbles. Two marbles are drawn from the bag. What is the probability that the second ball drawn will be red if replacement is NOT allowed?

A. 1/36
B. 1/12
C. 7/36
D. 2/9
E. 7/9
No math is needed here if we understand the following concept:
The probability of selecting X on the NTH pick is equal to the probability of selecting X on the FIRST pick.
Thus, P(red on the 2nd pick) = P(red on the first pick) = 2/9.

The correct answer is D.

Other problems that test this concept:
https://www.beatthegmat.com/beat-this-pr ... 85719.html
https://www.beatthegmat.com/probablity-ques-t60161.html
https://www.beatthegmat.com/manhattan-pr ... 89481.html (2 posts)
https://www.beatthegmat.com/a-box-contai ... 51368.html
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by prachi18oct » Mon Jul 27, 2015 6:05 am
Hi GMATGuruNY,

Will the probability of picking the second ball as red not change if the first ball picked is also red?
I did it as follows:-
P(second ball as red) = Prob of first red ball * prob of second red ball + prob of first non-red ball * prob of second ball as red
=> 2/9 * 1/8 + 7/9 * 2/8 = 16/72 = 2/9

Please advise.
Join the discussion

by GMATGuruNY » Mon Jul 27, 2015 6:37 am
prachi18oct wrote:I did it as follows:-
P(second ball as red) = Prob of first red ball * prob of second red ball + prob of first non-red ball * prob of second ball as red
=> 2/9 * 1/8 + 7/9 * 2/8 = 16/72 = 2/9
Your solution is perfect.
Will the probability of picking the second ball as red not change if the first ball picked is also red?
Yes.
As your solution illustrates, P(RR) ≠ P(NR):
P(RR) = (2/9)(1/8) = 2/72.
P(NR) = (7/9)(2/8) = 14/72.

However, the SUM of these probabilities is equal to the probability of selecting a red marble on the first pick:
2/72 + 14/72 = 16/72 = 2/9.

As your solution proves, if we account for ALL of the ways to select a red marble on the NTH pick, the result will be equal to the probability of selecting a red marble on the FIRST pick:
P(red marble on the nth pick) = P(red marble on the 1st pick) = 2/9.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by Brent@GMATPrepNow » Mon Jul 27, 2015 6:45 am
This question reminds me of my childhood, when my friends and I would sometimes "draw straws" to randomly select one person to do something (often either work, like getting wood for the fire, or dumb, like eating something that shouldn't be eaten).

So, someone would hold up n pieces of grass (for n people), and one of those pieces was very short. The person who selected the shortest piece was the one who had to perform the task.

There was always one guy who wanted to choose his piece last. His reasoning was that his chances of drawing the shortest piece were minimized since every person before him had a chance of drawing the short piece before it got to his turn.

The truth of the matter is that each of the n people had a 1/n chance of selecting the shortest piece, regardless of the order in which they selected.

The same applies to the original question here. 2 of the 9 balls are blue, so P(red ball selected second) = 2/9

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by nikhilgmat31 » Wed Jul 29, 2015 12:30 am
I did it in long way
selecting a White Ball at first & selecting red in second OR selecting red ball at first & selecting red in second OR selecting BLUE ball at first & selecting red in second

3/9 * 2/8 = 6/72
2/9 * 1/8 = 2/72
4/9 * 2/8 = 8/72

summing all 3 gives = 16/72 2/9

It is same as doing selecting non red & then selecting a red ball OR selecting RED ball and then selecting RED ball.

7/9 * 2/8 + 2/9 * 1/8

16/72 = 2/9
Join the discussion